OpenStudy (anonymous):

solve the following system of equations x + 3y = 5 (1) x = 4 - 3y (2) what is the solution of the system type and ordered pair

OpenStudy (anonymous):

dipankarstudy can you help me out solve the following system of equations x + 3y = 5 (1) x = 4 - 3y (2) what is the solution of the system type and ordered pair

OpenStudy (anonymous):

\[There\ is\ no\ solution\ for\ this\]

OpenStudy (anonymous):

can you show me how u arrived at the ans

OpenStudy (anonymous):

\[coz\ both\ x \ and\ y\ cancels\ out\]

OpenStudy (anonymous):

\[did\ u\ understand\]

OpenStudy (anonymous):

sorta